A father is x years old and his daughter is y years old. three times the difference in their ages is 57. In 2 years time the father will be twice as old as the daughter. How old are they now? I can't work out what the equation would be to work this out. I thought x-y=57/3 which is the age difference but beyond that I do not know how to approach this question. Thank you.
Answer by richwmiller(17219)

3*(f-d)=57
f-d=19
f=d+19
f+2=2*(d+2)
f+2=2d+4
f=2d+2
d+19=2d+2
17=d
f=36